Big points!!
Could someone please help me solve these?
I'm looking for formulas and necessary calculations so I can try myself.
Thank you!!
Ps the last word on #78 is "successfully."
media%2F660%2F660087d8-6f2d-45b5-ad75-99
Answer #1

73.
mass m = 516 kg

radius , r = 155m

speed v = 29*1609/3600 = 12.96 m/s

mu = 0.3

answer C) 13 m/s
------------------
74.
centripetal acceleration a = v^2/r = 1.08 m/s ^2

Answer:: c.1.08 m/s^2
---------------------
75.
force F= ma = 516*1.08 = 560 N

answer:: a.560 N
-------------------------

76.

maximum static friction = mu*mg = 0.3*516*9.81 = 1520 N

Answer::: a.1.52 x 10^3 N
-----------------------
77.
yes

------------------
78.
v =sqrt(0.3*516*9.81) = 21.3 m/s
Answer::: b. 21.3 m/s